How many grams of Oxygen are produced if 25g of potassium chlorate decompose?
1
Expert's answer
2019-12-05T08:39:33-0500

2KCIO3"\to" 2KCI+3O2

n(KCIO3)=(25(g)/122.4(g/mol)=2.05moles

n(O2)=2.05moles*3/2=3.075moles

m(O2)=3.075*32=98.4g